OpenStudy (mrs.ambrose614):

WILL MEDAL Which of the following statements best describes the graph of 4x − y = 1? It is a straight line joining the points (0, −1), (1, 3), and (−1, −5). It is a straight line joining the points (4, −3), (−1, 2), and (−4, 5). It is a curve joining the points (4, −1), (2, 3), and (4, 1). It is a curve joining the points (0, −1), (−1, −3), and (1, 5).

OpenStudy (sooobored):

which answer do you think it is and why?

OpenStudy (mrs.ambrose614):

I think it is c

OpenStudy (mrs.ambrose614):

But am I right

OpenStudy (sooobored):

WHYYYYY

OpenStudy (sooobored):

and no, you are incorrect

OpenStudy (sooobored):

you can check by substituting one of the points into the equation if we look at the first point of c (4,-1) substitution 4(4)- (-1) = 1 16 +1 =1 \(17 \ne 1\) hence c is incorrect

OpenStudy (mathmale):

Hint: "4x − y = 1?" is a LINEAR equation. Think: What does that mean? To what power are x and y raised? If you can answer this question correctly, you could immediately eliminate 2 of the 4 possible answer choices.